Mr.La Fleur pays $130 to rent a car for one week. He also pays $5 for every 100 miles that he drives. Write an equation and draw a graph which shows the relationship between the total cost of renting the car and the number of miles that are driven.
Answer by Mathtut(3670)

so we have total cost of let say C
C=130+5(m) where m is the total number of miles divided by 100. 130 is a fixed cost that doesnt change no matter how many miles are driven.
so make a graph where y axis is the cost and the x axis m(number of miles/100)
if miles driven is say 500 then m would be 500/100..so C=130+5(5)=155
so you would plot (5,155) on the graph(perhaps every mark on the graph could equal 5)
when miles driven is 1000, m would be 1000/100=10..so C=130+5(10)=180
(10,180) as you can see for every 500 miles you drive you add $25 to the cost.